if the vectors i ​− 2xj − 3yk and i + 3xj + 2yk are orthogonal to each other , then the locus of the point (x , y ) is 
 
  • hyperbola      
  • ellipse
  • parabola
  • circle 

Grade:12th pass

1 Answers

Mohd Mujtaba
6 years ago
Answer is "circle". If these two vector are orthogonal then their dot product is 0. So (i-2xj-3yk).(i+3xj+2yk)=0. On solving we get 6x^2+6y^2=1. Or x^2+y^2=1/(root6)^2. Which is equation of circle. Thank☺☺☺☺